OpenStudy (anonymous):

need help with solving systems of equations 4x+2y-6z=-38 5x-4y+z=-18 x+3y+7z=38

OpenStudy (anonymous):

so far I tried to combine all three and got 10x+y+2z=-18 now I'm not sure what to do

OpenStudy (jack1):

x = 38 - 3y - 7z (from 3rd equation) sub that into equation 2 5x-4y+z=-18

OpenStudy (anonymous):

you need to use the elimination method

OpenStudy (anonymous):

or substitution

OpenStudy (anonymous):

ok so substituting the third equation into the secon i'll get 5(38-3y-7z)-4y+z=-18

OpenStudy (anonymous):

190-15y-35z-4y+z=-18

OpenStudy (jack1):

5x - 4y + z =-18 5(38 - 3y - 7z) -4y + z = -18 will give you an answer purely in terms of y and z put in terms of z=.... sub that answer into eqn 1 along with your answerr from eqn 3 (x=...) wil give u an exact answer for y

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

ok right now I got 34z=208-19y

OpenStudy (jack1):

cool, so z=(208-19y)/34

OpenStudy (anonymous):

ok but how do I divide that not knowing what y is?

OpenStudy (jack1):

cool, so sub x = 38 - 3y - 7z into eqn 1 expand and add like terms then sub z=(208-19y)/34 you'll be left with y = answer

OpenStudy (anonymous):

I got z=(-190+10y)/34 and from equ.2 z=(208-19y)/34

OpenStudy (anonymous):

im confused.

OpenStudy (jack1):

yep, sub in x=... gives you: 10y + 34z = 190 and sub in z=... gives you 10y + 34((208-19y)/34) = 190 so 34/34 = 1/1 so left with: 10y + 208-19y = 190 = -9y = -18 y = -18/-9 = 2 therefore y = 2

OpenStudy (anonymous):

ooooooooooooooooo i think I got it

OpenStudy (jack1):

cools now use z=(208-19y)/34 to find z and x = 38 - 3y - 7z to find x done ;D

OpenStudy (jack1):

slaters, good luck

OpenStudy (anonymous):

x=-3, y=2, and z=5

OpenStudy (anonymous):

Thanks so much!!!!!!!!!!!!!!!!

OpenStudy (jack1):

perfect, props hey ;D